The specific heat of a substance varies with temperatureaccording to c = 0.20 +0.99T + 0.040T2,with T in °C andc in cal/g·K. Findthe energy required to raise the temperature of 7.4 g of thissubstance from 7.0°C to 25°C.
Answer #1
For convenience, let us put,
                    c = 0.20 + 0.99T +0.040T2 = a + bT+c.T2
For an infinitesimal rise in temperature, dT, theenergy required,
       dQ = m.c.dT = m( a +bT +c.T2).dT
          Q =Integral dQ
             =m.(a.T + 0.5bT2 + 0.333c.T3).
             evaluated between T = 7 and T = 25
Hence, Q = 7.4 g(0.2 x18 + 0.997x288 + 0.0407x5094)cal/g
               = 7.4 x 498 = 3685cal
Answer #2
dQ = m * c * dT     where Q is in caloriesand T specified to be Celsius
                                                                                  25
So Q = m * [.20 * T + .99 * T^2 / 2 + .04 * T^3 /3]       = (523 - 30) * 7 = 3450cal
                                                                                  7
E = 4.19 J / cal * 3450 cal = 14,460 J